# A Survey on Error-Bounded Lossy Compression for Scientific Datasets

### v1

@@ 3.5. Bit-Plane Coding (BPC)
Table 3. Bit-Plane Coding Methods Used in Different Lossy Compressors
@@
cuSZp (Laboratory, 2023) Processing Nonzero Blocks after Quantization+Lorenzo Absolute Error Bound
@@
3.6. Tucker Decomposition and HOSVD (SVD)
@@ 3.7. Decimation/Sampling (DS)
Decimation/sampling is commonly used by scientific applications to reduce the volumes of the simulation data to be stored on parallel file systems. In general, decimation means performing downsampling along the time dimension during the simulation: for example, saving the snapshot data to disks every $K$ time steps instead of saving all snapshots during the simulation. Many scientific simulation packages, such as Hardware/Hybrid Accelerated Cosmology Code (HACC) (Habib et al., 2013), EXAALT molecular dynamics simulation (exaalt, 2020), reverse time migration (RTM) (Bartana et al., 2015; rtm, [n. d.]), and Flash-X (fla, [n. d.]), allow users to save the snapshot data selectively over time. In comparison with decimation, the sampling strategy generally means performing downsampling in space for each snapshot dataset, which can also significantly reduce the data volumes. Liang et al. (Liang et al., [n. d.]) studied the pros and cons of different decimation/sampling-based compression strategies in both temporal and spatial dimension. Specifically, the authors pointed out that a decimation/sampling method can have extremely high speed in the compression stage, but it may suffer from substantial decompression cost and also low reconstructed data quality compared with traditional error-bounded lossy compressors such as SZ (Tao et al., 2017c; Liang et al., 2018b). Compressed sensing (CS) is another typical lossy compression method that leverages the sampling strategy. In general, CS is used where the compression is required to be very fast (e.g., in online compression) while decompression is not that important and can be performed offline. CS can be very fast because it just needs to sample the dataset with a certain randomness. To reconstruct the data, however, CS needs to solve an underdetermined linear system, which could be very expensive.
3.8. Filtering (FTR)
@@
The filtering technique aims to remove insignificant values or ignore the insignificant changes of the data, which can then significantly reduce the data size. In general, the significance of the data is determined by the impact of the data being processed on the final reconstructed data quality. The filtering technique has been widely used in many existing error-bounded lossy compressors, such as (cu)SZx, cuSZp, and SPERR[] and the specific filtering methods often appear in different forms. In the following, we describe two forms of filtering commonly used in lossy compressors.
@@
  • Data Folding. Data folding aims to replace (“fold”) a set of data with one single value, provided that the variation of these data can be ignored. The error-bounded compressor SZx is a good example. SZx splits the whole dataset into many fixed-length consecutive 1D blocks. If all the data values in a block are close to each other such that the value interval range of the block is lower than or equal to twice the user-required error bound, then the mean of the min and max in this block can be used to replace all values in the block. Such blocks are called “constant blocks” in SZx. Similarly, cuSZp (Laboratory, 2023) also splits the whole dataset into many blocks, and each block performs uses quantization and Lorenzo prediction to decorrelate the data. Most of data values then tend to be very close to 0, and the blocks with all zeros would be just represented by a 1-byte mark. This is essentially a type of data folding.
@@ 3.9. Lossless Encoding (LE)
RunLength Encoding (RE) cuSZ+, TTHRESH Reduce Repeated Symbols (Tian et al., 2021; Ballester-Ripoll et al., 2019)
@@
3.10. Deep Neural Network (DNN)
@@
Since neural-network-based compression has been well developed and practicalized for natural images (Hu et al., 2021; Mishra et al., 2022; Jamil et al., 2023) and videos (Bidwe et al., 2022), several initial attempts have also been made to leverage neural networks for the lossy compression of scientific data. In neural-network-based scientific lossy compressors, the neural networks can serve as both data encoders (Choi et al., [n. d.]; Lu et al., 2021; Liu et al., 2021e; Glaws et al., 2020; Hayne et al., 2021; Huang et al., 2021, 2023d; Liu et al., 2021a) and data predictors (Huang and Hoefler, 2022; Han and Wang, 2022; Han et al., 2023; Liu et al., 2023b) and can also be offline-pretrained by preacquired datasets (Choi et al., [n. d.]; Liu et al., 2021e; Glaws et al., 2020; Hayne et al., 2021; Huang et al., 2021, 2023d; Liu et al., 2023b, 2021a) or be online-trained by input data (Lu et al., 2021; Huang and Hoefler, 2022; Han and Wang, 2022; Han et al., 2023). For example, AE-SZ (Liu et al., 2021a) encodes the input data with a pretrained convolutional Sliced-Wasserstein Autoencoder (SWAE), and SRNN-SZ applies a pretrained Hybrid Attention Transformer (HAT) as an interpolation-like data predictor. CoordNet (Han and Wang, 2022) and KD-INR (Han et al., 2023) are examples of leveraging predictive neural networks with data coordination information in scientific compression, in which the networks are online trained by the input data before the data prediction process. Neural network-based compressors with online-trained networks can achieve much better compression ratios and/or distortions than offline-trained networks but suffer from lower throughputs due to the requirement of training for each separate input. Neural-network-based compressors with offline-trained networks are free from per-input training but also need to address the challenge of collecting trustworthy training datasets. No matter which scheme is used, neural-network-based scientific lossy compressors still must overcome the limitation of low-speed neural networks, presenting a better balance of quality and performance to fit the practical usage in high-performance scientific computing systems.
4. General-Purpose Error-Bounded Lossy Compressors
In Table 5 we summarize the compression pipelines used by 35 error-bounded lossy compressors. We describe 12 representative state-of-the-art lossy compressors in detail in the following text of this section.
Table 5. 36 Error-Bounded Compressors Each with Various Pipelines, Domains, Targeted Devices, and Datasets
Compressor Year Compression Pipeline Domain Device Dataset References
FPZIP 2006 PDP+BPC Generic CPU Structured (Lindstrom et al., 2017)
ISABELA 2013 Sorting+DS+Bspline Generic CPU Structured (Lakshminarasimhan et al., 2011, 2013)
ZFP 2014 PDT+OWT+BPC Generic CPU/GPU Structured (Lindstrom, 2014; cuZFP, 2020)
SZ0.1 2016 PDP+LE Generic CPU Structured (Di and Cappello, 2016)
Bitgrooming 2016 BPC+LE Generic CPU Structured (Zender, 2016)
SZ1.4 2017 PDP+QT+LE Generic CPU Structured (Tao et al., 2017c)
SZ2 2018 PDP+QT+LE Generic CPU Structured (Liang et al., 2018b)
Pastri-SZ 2018 PDP+QT+LE Quantum Chemistry CPU Structured (Gok et al., 2018)
MGARD 2018 PDP+QT+LE Generic CPU (Un)structured (Ainsworth et al., 2018, 2020)
Digitrounding 2019 BPC+LE Generic CPU Structured (dig, [n. d.])
DCTZ 2019 OWT+QT+LE Generic CPU Structured (Zhang et al., 2019)
GhostSZ 2019 PDP+QT+LE Generic FPGA Structured (Xiong et al., 2019)
TTHRESH 2019 SVD+DS Generic CPU Structured (Ballester-Ripoll et al., 2019)
ZFP-V 2019 PDT+OWT+BPC Generic FPGA Structured (Sun and Jun, 2019)
DeepSZ 2019 PDP+QT+LE DNN model CPU Structured (Jin et al., 2019)
SZauto 2020 PDP+QT+LE Generic CPU Structured (Zhao et al., 2020b)
cpSZ 2020 PDP+QT+LE Critical Points CPU (Un)structured (Liang et al., 2020, 2023)
in Vector Fields
waveSZ 2020 QT+PDP+LE Generic FPGA Structured (Tian et al., 2020a)
cuSZ 2021 QT+PDP+HE+RE Generic GPU Structured (Tian et al., 2020b, 2021)
SZ 3 2021 PDP+QT+LE Generic CPU Structured (Zhao et al., 2021; Liang et al., 2021)
AESZ 2021 DNN+PDP+QT+LE Generic CPU Structured (Liu et al., 2021a, e)
GPU-MGARD 2021 PDP+QT+HE+RE Generic GPU Structured (Chen et al., 2021; Gong et al., 2023)
DE-ZFP 2022 PDT+OWT+BPC Generic FPGA Structured (Habboush et al., 2022)
QOZ 2022 PDP+QT+LE Generic CPU Structured (Liu et al., 2022b)
MDZ 2022 PDP+QT+LE Molecular Dynamics CPU Structured (Zhao et al., 2022)
SZx 2022 FTR+BPC Generic CPU/GPU Structured (Yu et al., 2022)
SPERR 2023 DWT+QT+LE Generic CPU Structured (Li et al., 2023)
FAZ 2023 DWT+PDP+QT+LE Generic CPU Structured (Liu et al., 2023c)
FP-GPU 2023 CE+QT+PDP+PE Generic GPU Structured (Zhang et al., 2023)
cuSZp 2023 CE+QT+PDP+PE Generic CPU/GPU Structured (Huang et al., 2023a; Laboratory, 2023)
Roibin-SZ 2023 PDP+QT+LE Light-source Data CPU/GPU Structured (Underwood et al., 2023)
AMR-SZ 2023 PDP+QT+LE Generic CPU Unstructured (Wang et al., 2022, 2024, 2023b, 2023a)
SRNN-SZ 2023 DNN+PDP+QT+LE Generic CPU Structured (Liu et al., 2023b)
topoSZ 2023 PDP+QT+LE Contour Tree in CPU Structured (Yan et al., cess)
Scalar Fields
FedSZ 2023 PDP+QT+LE Federated Learning CPU Structured (Wilkins et al., 2023)
CliZ 2024 PDP+QT+LE Climate Research CPU Structured (Jian et al., 2024)
4.1. SZ

### v2

@@ 4.5 Bit-Plane Coding (BPC)
Table 4. Bit-Plane Coding Methods Used in Lossy Compressors: bit-plane coding is used in many compressors for various purposes, also lreadig to diverse error control modes.
@@
cuSZp (Huang et al., 2023a) Processing Nonzero Blocks after Quantization+Lorenzo Absolute Error Bound
@@
4.6 Tucker Decomposition and HOSVD (SVD)
@@ 4.7 Decimation/Sampling (DS)
Decimation/sampling is commonly used by scientific applications to reduce the volumes of the simulation data to be stored on parallel file systems. In general, decimation means performing downsampling along the time dimension during the simulation: for example, saving the snapshot data to disks every $K$ time steps instead of saving all snapshots during the simulation. Many scientific simulation packages, such as Hardware/Hybrid Accelerated Cosmology Code (HACC) (Habib et al., 2013), EXAALT molecular dynamics simulation (exa, 2020), reverse time migration (RTM) (Bartana and et al., 2015; rtm, [n. d.]), and Flash-X (fla, [n. d.]), allow users to save the snapshot data selectively over time. In comparison with decimation, the sampling strategy generally means performing downsampling in space for each snapshot dataset, which can also significantly reduce the data volumes. Liang et al. (Liang et al., [n. d.]) studied the pros and cons of different decimation/sampling-based compression strategies in both temporal and spatial dimension. Specifically, the authors pointed out that a decimation/sampling method can have extremely high speed in the compression stage, but it may suffer from substantial decompression cost and also low reconstructed data quality compared with traditional error-bounded lossy compressors such as SZ (Tao et al., 2017c; Liang et al., 2018b). Compressed sensing (CS) is another typical lossy compression method that leverages the sampling strategy. In general, CS is used where the compression is required to be very fast (e.g., in online compression) while decompression is not that important and can be performed offline. CS can be very fast because it just needs to sample the dataset with a certain randomness. To reconstruct the data, however, CS needs to solve an underdetermined linear system, which could be very expensive.
4.8 Filtering (FTR)
@@
The filtering technique aims to remove insignificant values or ignore the insignificant changes of the data, which can then significantly reduce the data size. In general, the significance of the data is determined by the impact of the data being processed on the final reconstructed data quality. The filtering technique has been widely used in many existing error-bounded lossy compressors, such as (cu)SZx, cuSZp, and SPERR and the specific filtering methods often appear in different forms. In the following, we describe two forms of filtering commonly used in lossy compressors.
@@
  • Data Folding. Data folding aims to replace (“fold”) a set of data with one single value, provided that the variation of these data can be ignored. The error-bounded compressor SZx is a good example. SZx splits the whole dataset into many fixed-length consecutive 1D blocks. If all the data values in a block are close to each other such that the value interval range of the block is lower than or equal to twice the user-required error bound, then the mean of the min and max in this block can be used to replace all values in the block. Such blocks are called “constant blocks” in SZx. Similarly, cuSZp (Huang et al., 2023a) also splits the whole dataset into many blocks, and each block performs uses quantization and Lorenzo prediction to decorrelate the data. Most of data values then tend to be very close to 0, and the blocks with all zeros would be just represented by a 1-byte mark. This is essentially a type of data folding.
@@ 4.9 Lossless Encoding (LE)
RunLength Encoding (RE) cuSZ+, TTHRESH Reduce Repeated Symbols (Tian and et al., 2021; Ballester-Ripoll et al., 2019)
@@
4.10 Deep Neural Network (DNN)
Figure 5. Illustration of Autoencoder (AE): Encoder and decoder are two models trained with datasets, and latent vector represents the space in compressed format.
@@ 4.10 Deep Neural Network (DNN)
Since neural-network-based compression has been well developed and practicalized for natural images (Hu et al., 2021; Mishra et al., 2022; Jamil et al., 2023) and videos (Bidwe et al., 2022), several initial attempts have also been made to leverage neural networks for the lossy compression of scientific data. In neural-network-based scientific lossy compressors, the neural networks can serve as both data encoders (Choi et al., [n. d.]; Lu et al., 2021; Liu et al., 2021d; Glaws et al., 2020; Hayne et al., 2021; Huang et al., 2021, 2023b; Liu et al., 2021a) and data predictors (Huang and Hoefler, 2022; Han and Wang, 2022; Han et al., 2023; Liu et al., 2023b) and can also be offline-pre trained by pre-acquired datasets (Choi et al., [n. d.]; Liu et al., 2021d; Glaws et al., 2020; Hayne et al., 2021; Huang et al., 2021, 2023b; Liu et al., 2023b, 2021a) or be online-trained by input data (Lu et al., 2021; Huang and Hoefler, 2022; Han and Wang, 2022; Han et al., 2023). Auto-encoder (AE/VAE) can effectively reduce the dimensions or compress original data into a latent vector, as illustrated in Figure 5. Simply using auto-encoder (AE/VAE), however, may introduce unaccepted errors (Liu et al., 2021d), so many AE-based compressors leveraged AE as a predictor, followed by quantization and lossless encoding. For example, AE-SZ (Liu et al., 2021a) encodes the input data with a pretrained convolutional Sliced-Wasserstein Autoencoder (SWAE), and SRNN-SZ applies a pre trained Hybrid Attention Transformer (HAT) as an interpolation-like data predictor. These two compressors are typical examples with offline-trained networks. Another example (HH-NN) will be described in Section 6.4. CoordNet (Han and Wang, 2022) and KD-INR (Han et al., 2023) are two compressors in which the networks are online trained by the input data before the data prediction process. Neural network-based compressors with online-trained networks can achieve much better compression ratios and/or distortions than offline-trained networks for general datsaets but suffer lower throughputs due to the requirement of training for each separate input.
5. General-Purpose Lossy Compressors for Scientific Data
In this section, we describe general-purpose lossy compressors, each of which projects a specific compression pipeline composed of a series of steps or techniques/modules. The selection of lossy compressorsin our study is based on the following criteria.
  • These compressors are designed for scientific datasets generated by HPC applications or advanced X-ray facilities such as APS (Fornek, 2017).
  • These compressors provide certain compression error controls, in the form of error bound (Di and Cappello, 2016; Tao et al., 2017c), bit-truncation precision (Lindstrom et al., 2017; Lindstrom, 2014), or control of impact to the posthoc analysis.
Table 6 summarizes 46 lossy compressors, each with a distinct design or compression pipeline. In the last two columns of Table 6, we present an examination of the performance (compression/decompression speed) and efficiency (compression ratio) for all the listed compressors, based on published results. Performance and compression ratio are categorized into five levels, with more stars indicating higher levels. The levels are also determined considering the device and the comparison among other works. For example, SPERR, HH-NN and FAZ have the top level of compression ratio (*****) because their compression ratios can be 1-2 orders of magnitude higher than that of other state-of-the-art compressors like SZ and ZFP. CuSZp is ranked as the top level of the performance because it features a very high end-to-end compression/decompression throughput on GPU (200-400 GB/s on a single A100 GPU), which is significantly higher than other GPU compressors do (Huang et al., 2024a). Note that this rating is a rough estimation based on the literature: e.g., the compression ratio level * means 1$\sim$5 while the level ***** means one thousand or even higher at the similar level of reconstructed data quality (in terms of PSNR or visualization).
Table 6. Summary of General-purpose Lossy Compressors
Lossy Compressors Year Model Compress. Pipeline Device Type QoI Performance Compression Ratio FPZIP (Lindstrom et al., 2017) 2006 Pred. PDP+BPC CPU S N *** (Di and Cappello, 2016) ** (Liang et al., 2018b; Di and Cappello, 2016) ISABELA (Lakshminarasimhan et al., 2011, 2013) 2013 Deci. Sorting+DS+Bspline CPU S N * (Di et al., 2018) * (Di et al., 2018) ZFP (Lindstrom, 2014; cuZFP, 2020) 2014 Trans. DT+OWT+BPC CPU/GPU S N **** (Tao et al., 2017c; Liang et al., 2018b) ****(Tao et al., 2017c; Liang et al., 2018b) NUMARCK (Chen et al., 2014) 2014 Pred. PDP+QT CPU S N - ** (Di and Cappello, 2016) SSEM (Sasaki et al., 2015) 2015 Trans. DWT+QT+FTR+LE CPU S N - ** (Di and Cappello, 2016) SZ0.1 (Di and Cappello, 2016) 2016 Pred. PDP+LE CPU S N *** (Di and Cappello, 2016) *** (Di and Cappello, 2016) Bitgrooming (Zender, 2016) 2016 BitM. BPC+LE CPU S N - * (Underwood et al., 2022) SZ1.4 (Tao et al., 2017c) 2017 Pred. PDP+QT+LE CPU S N *** (Tao et al., 2017c; Liang et al., 2018b) *** (Tao et al., 2017c; Liang et al., 2018b) SZ2 (Liang et al., 2018b) 2018 Pred. PDP+QT+LE CPU S N **** (Liang et al., 2018b; Zhao et al., 2021) **** (Liang et al., 2018b; Zhao et al., 2021) MGARD (Ainsworth et al., 2018, 2020) 2018 Pred. PDP+QT+LE CPU U&S Y ** (Liang et al., 2021b) *** (Liang et al., 2021b) SpaioTmpDeci. (Liang et al., [n. d.]) 2018 Decim. DS CPU S N **** (Liang et al., [n. d.]) * (Liang et al., [n. d.]) Digitrounding (dig, [n. d.]) 2019 BitM. BPC+LE CPU S N - * (Underwood et al., 2022) DCTZ (Zhang et al., 2019b) 2019 Trans. OWT+QT+LE CPU S N *** (Zhang et al., 2019b) *** (Zhang et al., 2019b) GhostSZ (Xiong et al., 2019) 2019 Pred. PDP+QT+LE FPGA S N *** (Xiong et al., 2019) **** (Xiong et al., 2019) TTHRESH (Ballester-Ripoll et al., 2019) 2019 HOSVD SVD+BPC+LE CPU S N * (Ballester-Ripoll et al., 2019; Liu et al., 2024; Baert and Vannieuwenhoven, 2023) ***** (Baert and Vannieuwenhoven, 2023) ZFP-V (Sun and Jun, 2019) 2019 Trans. DT+OWT+BPC+LE FPGA S N **** (Sun and Jun, 2019) **** (Sun and Jun, 2019) SZauto (Zhao et al., 2020b) 2020 Pred. PDP+QT+LE CPU S N *** (Zhao et al., 2020b) **** (Zhao et al., 2020b) waveSZ (Tian and et al., 2020) 2020 Pred. QT+PDP+LE FPGA S N *** (Tian and et al., 2020) **** (Tian and et al., 2020) TuckerMPI (Ballard et al., 2020) 2020 HOSVD SVD CPU S N ** (Baert and Vannieuwenhoven, 2023) ** (Baert and Vannieuwenhoven, 2023) zMesh(Luo et al., 2021) 2021 Pred. DT CPU U N *** (Luo et al., 2021) **** (Luo et al., 2021) TEZip (Roy and et al., 2021) 2021 DeepL. DNN+LE CPU S N **** (Roy and et al., 2021) **** (Roy and et al., 2021) cuSZ (Tian et al., 2020; Tian and et al., 2021) 2021 Pred. QT+PDP+LE GPU S N ** (Huang et al., 2023a) *** (Huang et al., 2023a, 2024a; Liu and et al., 2024) MGARD+ (Liang et al., 2021b) 2021 Pred. PDP+QT+LE CPU S N *** (Liang et al., 2021b) **** (Liang et al., 2021b) SZ3 (Zhao et al., 2021; Liang et al., 2021a) 2021 Pred. PDP+QT+LE CPU S N *** (Zhao et al., 2021) **** (Zhao et al., 2021; Liu et al., 2024, 2023c) AESZ (Liu et al., 2021a, d) 2021 DeepL. DNN+PDP+QT+LE CPU S N ** (Liu et al., 2021a) **** (Liu et al., 2021a) CAE (Liu et al., 2021d) 2021 DeepL. DNN+PDP+QT+LE CPU S N ** (Liu et al., 2021d) **** (Liu et al., 2021d) MGARD-GPU(Chen et al., 2021; Gong et al., 2023) 2021 Pred. PDP+QT+LE GPU S N *** (Gong et al., 2023) *** (Gong et al., 2023) DE-ZFP(Habboush et al., 2022) 2022 Trans. DT+OWT+BPC+LE FPGA S N **** (Habboush et al., 2022) **** (Habboush et al., 2022) QoZ (Liu et al., 2022b) 2022 Pred. PDP+QT+LE CPU S N *** (Liu et al., 2022b, 2024) **** (Liu et al., 2022b, 2024) SZx (Yu et al., 2022) 2022 BitM. FTR+BPC CPU S N ***** (Yu et al., 2022) ** (Yu et al., 2022) cuSZx (Yu et al., 2022) 2022 BitM. FTR+BPC GPU S N ** (Huang et al., 2023a) ** (Huang et al., 2023a) Jiao et al. (Jiao et al., 2022) 2022 Pred. DT+PDP+QT+LE CPU S Y ** (Jiao et al., 2022) *** (Jiao et al., 2022) MGARD-Lambda (mga, [n. d.]) 2022 Pred. PDP+QT+LE CPU S Y - - ATC (Baert and Vannieuwenhoven, 2023) 2023 HOSVD SVD+QT+BPC+LE CPU S N ** (Baert and Vannieuwenhoven, 2023) ***** (Baert and Vannieuwenhoven, 2023) ZFP-X (Lu et al., 2023) 2023 Trans. DT+OWT+BPC+LE CPU S N **** (Lu et al., 2023) **** (Lu et al., 2023) SPERR (Li et al., 2023) 2023 Trans. DWT+QT+LE CPU S N ** (Li et al., 2023; Liu et al., 2023c) ***** (Li et al., 2023; Liu et al., 2023c) HH-NN (Huang and Hoefler, 2023) 2023 DeepL. DS+DNN+QT GPU S N * (Huang and Hoefler, 2023) ***** (Huang and Hoefler, 2023) FAZ (Liu et al., 2023c) 2023 Pred. DWT+PDP+QT+LE CPU S N ** (Liu et al., 2023c) ***** (Liu et al., 2023c) FZ-GPU (Zhang et al., 2023) 2023 Pred. QT+PDP+LE GPU S N **** (Zhang et al., 2023) ** (Zhang et al., 2023) cuSZp (Huang et al., 2023a, 2024a) 2023 Pred. QT+PDP+LE GPU S N ***** (Huang et al., 2023a, 2024a) ** (Huang et al., 2023a, 2024a) AMR-Comp (Wang et al., 2022, 2024b, 2023b, 2023a, 2024a) 2023 Pred. PDP+QT+LE CPU U N *** (Wang et al., 2024b) **** (Wang et al., 2024b) SRNN-SZ (Liu et al., 2023b) 2023 Pred. DNN+PDP+QT+LE CPU S N * (Liu et al., 2023b) ***** (Liu et al., 2023b) SZ_ADT (Lu and et al., 2023) 2023 Pred. PDP+QT+LE CPU S N *** (Lu and et al., 2023) **** (Lu and et al., 2023) HPEZ(QOZ2) (Liu et al., 2024) 2024 Pred. PDP+QT+LE CPU S N *** (Liu et al., 2024) **** (Liu et al., 2024) HAE (Le et al., 2024) 2024 DeepL. DT+DNN+QT+FTR GPU S N - **** (Le et al., 2024) SZp (szp, [n. d.]; Huang and et al., 2024a) 2024 Pred. QT+PDP+LE GPU S N ***** (Huang and et al., 2024a) ** (Huang and et al., 2024a)
Footnote: For the type column: S means Structured, U means Unstructured. Perf&Ratio column gives a rough ranking based on literature: more ‘*’ means higher throughput or ratio). The QoI column indicates whether the compressor supports preservation of Quantity of Interest (QoI). ‘-’ means no published sources to indicate its performance or compression ratio. 31 compressors, highlighted in bold font, are described more or less in the following text.
Table 7. Exemplifying multiple applications are tested by general-purpose compressors based on (Liu et al., 2024; Li et al., 2023; Lindstrom, 2014; Ballester-Ripoll et al., 2019) (MD: Molecular Dynamics, COM: Combustion, COS: Cosmology, CL: Climate, CH: Chesmitry, TU: Turbulence, GE: Geology, FU: Fusion.)
Compressor Application Datasets Tested
SZ1/2/3 MD, COM, COS, CL, CH, TU, GE, FU, etc.
ZFP COM, COS, CL, CH, GE, TU, etc.
MGARD COM, COS, CL, CH, GE, TU, etc.
TTHRESH COM, CL, GE, TU, etc.
SPERR COM, COS, CL, CH, GE, TU, etc.
In the following text, we use some representative state-of-the-art compressors to describe the compression design/pipelines and unique features. The lossy compressors selected here for further description represent typical models/families in lossy compression and they are featured in the “Compression for Scientific Data” tutorial at SC24 (Cappello et al., 2024). Many of them have been widely used in the community and frequently evaluated in lossy compression and compressibility analysis studies, including (Zhao et al., 2020a; Lu et al., 2018; Liang et al., 2018b; Underwood et al., 2023a; Rahman et al., 2024; Rahman and et al., 2023; Nguyen et al., 2024; Liu and et al., 2024). We exemplify some example scientific applications used to evaluate different compressors (identified by existing publications (Liu et al., 2024; Li et al., 2023; Lindstrom, 2014; Ballester-Ripoll et al., 2019)), in Table 7.